Jessie is four years older than Lea. Eight years ago, he was twice as old as Lea. Find their present ages. Solve using two variables.
Answer by nerdybill(7384)

Jessie is four years older than Lea. Eight years ago, he was twice as old as Lea. Find their present ages. Solve using two variables.
.
Let J = Jessie's present age
and L = Lea's present age
.
Since we have two unknowns, we'll need two equations.
From, "Jessie is four years older than Lea." we get equation 1:
J = L+4
.
And from, "Eight years ago, he was twice as old as Lea." we get equation 2:
J-8 = 2(L-8)
.
Using the definition of "J" from equation 1:
J = L+4
Substitute it into equation 2:
J-8 = 2L
(L+4)-8 = 2(L-8)
L+4-8 = 2L-16
L-4 = 2L-16
12 years = L (Lea's present age)
.
Plug the above into equation 1 and solve for J:
J = L+4
J = 12+4
J = 16 years (Jessie's present age)